### v1

# Optimality of Frequency Moment Estimation

Abstract
Estimating the second frequency moment of a stream up to $(1\pm\varepsilon)$ multiplicative error requires at most $O(\log n/\varepsilon^{2})$ bits of space, due to a seminal result of Alon, Matias, and Szegedy. It is also known that at least $\Omega(\log n+1/\varepsilon^{2})$ space is needed. We prove an optimal lower bound of $\Omega\left(\log\left(n\varepsilon^{2}\right)/\varepsilon^{2}\right)$ for all $\varepsilon=\Omega(1/\sqrt{n})$. Note that when $\varepsilon>n^{-1/2+c}$, where $c>0$, our lower bound matches the classic upper bound of AMS. For smaller values of $\varepsilon$ we also introduce a revised algorithm that improves the classic AMS bound and matches our lower bound. Our lower bound holds also for the more general problem of $p$-th frequency moment estimation for the range of $p\in(1,2]$, giving a tight bound in the only remaining range to settle the optimal space complexity of estimating frequency moments.
1 Introduction
An extensive body of literature is devoted to the streaming model of computation, which is important for the analysis of massive datasets and in network traffic monitoring. A central problem in this model is the frequency moment estimation problem: Elements from a universe $U$ are given to the algorithm one-by-one, defining a vector of frequencies — that is, $f_{x}\in\mathbb{N}$ is the number of times the element $x\in U$ appeared in the stream; Finally, the algorithm has to return, with good probability, a $(1\pm\varepsilon)$-estimation of $F_{p}:=\sum_{x\in U}f_{x}^{p}$ — the $p$-th frequency moment of the stream. We generally denote the length of the stream by $n$ and assume that $|U|=\text{poly}(n)$. The main complexity parameter studied in this model is how much space is needed for the algorithm to succeed. The study of both the streaming model and of frequency moment estimation in it was initiated in the seminal 1996 work of Alon, Matias, and Szegedy [AMS96].
The case of $p=2$, or second moment estimation, is of particular importance. It is often called the repeat rate or surprise index, and is used in various tasks such as database query optimization [AGMS99], network traffic anomaly detection [KSZC03], approximate histogram maintenance [GGI+02] and more. Other moments of particular interest are $p=1$, corresponding to the approximate counting problem [Mor78, NY22], and $p=0$, corresponding to the distinct elements problem [FM85, IW03, KNW10b]. Among these special cases, only the space complexity of the first remains not fully understood. The original algorithm for $F_{2}$-estimation given by Alon, Matias, and Szegedy uses $O(\log n/\varepsilon^{2})$ bits of space; while the highest known lower bound due to Woodruff in 2004 [Woo04] is $\Omega(\log n+1/\varepsilon^{2})$ — leaving up to a quadratic gap between the upper and lower bounds for certain choices of $\varepsilon$.
While $F_{p}$-estimation for $p\leq 2$ uses amount of space that is only logarithmic in the length of the stream, it was shown that for $p>2$ at least $\Omega(n^{1-2/p}/\text{poly}(\varepsilon))$ space is needed [BYJKS04, CKS03b] — which is polynomial in the stream’s length. A long list of works [IW05, BGKS06, MW10, AKO11, BO10, And17, Gan11b, WZ12, Gan11a, LW13] resulted in a nearly-tight bound of $\tilde{\Theta}\left(n^{1-2/p}/\varepsilon^{2}\right)$ for $F_{p}$-estimation for every $p>2$ (not necessarily an integer) and $\varepsilon$, for some ranges of parameters the bounds are tight — in others there is a gap between the bounds that is poly-logarithmic in the bound itself.
For $p\leq 2$ the space complexity is not as well understood. Woodruff [Woo04] showed a lower bound of $\Omega(\log n+1/\varepsilon^{2})$ for every $p\neq 1$, this is optimal in terms of $\varepsilon$ alone and is also known to be optimal for the distinct elements problem (that is, $p=0$). For the special case of approximate counting (that is, $p=1$), a tight bound of $\Theta(\log\log n+\log\varepsilon^{-1})$ is known [NY22]. For the range of $p\in[0,1)$, the upper bound of AMS was improved by Jayaram and Woodruff who presented a nearly-tight $\tilde{O}\left(\log n+1/\varepsilon^{2}\right)$ bound in that range [JW19]. This leaves $p\in(1,2]$ as the last remaining range within no nearly-tight bounds are known.
For certain generalizations more is known: When the stream is randomly shuffled and given in random order, then in the range $p\in(1,2)$ (excluding $p=2$) [BVWY18] showed an improved upper bound of $\tilde{O}\left(\log n+1/\varepsilon^{2}\right)$. When updates are allowed in the stream, that is, elements can also be deleted and not only added to it, then [KNW10a] showed that $\Theta(\log n/\varepsilon^{2})$ is optimal for $p\leq 2$.
In this work, we settle the space complexity of frequency moment estimation in the entire remaining range of $p\in(1,2]$, including the special case of second frequency moment estimation. For $p=2$, we show that the AMS bound is essentially tight.
Theorem.
Let $\mathcal{A}$ be a streaming algorithm that gives an $\left(1\pm\varepsilon\right)$ multiplicative approximation to the $F_{2}$ of its input stream and succeeds with probability $\geq 2/3$, for some $\varepsilon=\Omega(1/\sqrt{n})$. Then, the space used by $\mathcal{A}$ is $\Omega\left(\log\left(\varepsilon^{2}n\right)/\varepsilon^{2}\right)$.
Note that the range $\varepsilon<1/\sqrt{n}$ is less interesting as $O\left(\min\{n\log n,|U|\}\right)$ space suffices for exactly maintaining the vector of frequencies. We observe that in the range where $\varepsilon$ is very close to $1/\sqrt{n}$ our lower bound is (slightly) lower than the AMS upper bound, we show that this is inherent by introducing a modification of the AMS algorithm that matches our lower bound in this range.
Theorem.
For $\varepsilon=\Omega(1/\sqrt{n})$, we can get a $(1\pm\varepsilon)$-approximation of the $F_{2}$ of a stream of length $n$ using $O\left(\log\left(\varepsilon^{2}n\right)/\varepsilon^{2}\right)$ space with success probability $>2/3$.
$10\leq p<$ 1 $p=1$ $1<p\leq 2$ $p>2$
$\tilde{\Theta}(\log n+1/\varepsilon^{2})$ $\Theta(\log\log n+\log\varepsilon^{-1})$ $\Theta(\log n/\varepsilon^{2})$ $\tilde{\Theta}(n^{1-2/p}/\varepsilon^{2})$
Figure 1: Space complexity of $F_{p}$-estimation.
We also extend our lower bound to the range $p\in(1,2]$, which settles the space complexity of $F_{p}$-estimation for all values of $p$. See Figure 1 for a summary of the space complexity of $F_{p}$-estimation for all $p\geq 0$.
Theorem.
Fix $p\in(1,2]$. Let $\mathcal{A}$ be a streaming algorithm that gives an $\left(1\pm\varepsilon\right)$ approximation to the $F_{p}$ of its input stream, for some $\varepsilon\in\left(\Omega\left(n^{-1/p}\right),{4^{-1/(p-1)}}\right)$, and succeeds with probability $\geq 2/3$. Then, the space used by $\mathcal{A}$ is $\Omega\left(\log\left(\varepsilon^{1/p}n\right)/\varepsilon^{2}\right)$.
Most of the lower bounds for streaming problems are based on reductions from communication complexity. In [JW19], a natural barrier to prove a better than $\tilde{\Omega}(1/\varepsilon^{2})$ lower bound was shown: Even in a very strong model of communication, $O\left(1/\varepsilon^{2}\cdot\left(\log\log n+\log d+\log\varepsilon^{-1}\right)\right)$ bits of communication suffice for the players to correctly produce a $F_{p}$ estimation, where $d$ is the diameter of the communication graph. This means that problems who reduce to $F_{p}$-estimation have a too low communication complexity to improve the existing lower bounds. To overcome this natural barrier, we present a new type of a direct sum theorem that takes place at the level of the streaming algorithm rather than the level of the communication model — informally, we pack many instances of problems with communication complexity $\Theta(1/\varepsilon^{2})$ into a single stream, and then directly show that a successful streaming algorithm must solve them all. In Section 2 we give a detailed high-level overview of our proofs. The lower bound is presented in Section 4 and Section 5, and then extended from $p=2$ to $p\in(1,2]$ in Section 5.3. The improved upper bound is presented in Section 6. We conclude and present remaining open problems in Section 7.
2 High-Level Overview
In this Section we give a high-level overview of the components used in our lower and upper bounds. The lower bound is then proven in Section 4 and Section 5, and the upper bound in Section 6.
2.1 Exam Communication Model
As in many other streaming lower bounds, our initial building blocks are reductions from communication complexity problems. However, instead of using the classic communication model, we consider a slight variant which we call the exam model. In the classic numbers-in-hand communication model, $t$-players receive different parts $x_{1},x_{2},\ldots,x_{t}$ of an input and have to compute some function $f(x_{1},\ldots,x_{n})$ of the entire input by communicating with each other. In our model, we introduce an additional player, a referee, who has an additional input $y$ which we think of as a question about the players’ inputs $\{x_{i}\}$. The players still receive their parts of the input and are allowed to communicate with each other, but not with the referee; finally, a player sends a single message to the referee, and the referee then has to compute some function $g(x_{1},\ldots,x_{t},y)$ of both the input and her “secret” question $y$.
In the classic Disjointness problem, each of the $t$ players receives a set $S_{i}$ and they have to determine whether or not all of their input sets are pairwise disjoint. It is known that solving this problem remains hard even under the promise that the input sets are either all disjoint or contain a unique element appearing in all input sets and do not intersect further. We introduce a variant we call Exam Disjointness in which each of the players still receives a set $S_{i}$, and the referee receives a single element $y$ from the universe; based on the message she receives the referee needs to decide whether $y$ appears in all input sets $S_{i}$. In other words, instead of determining whether there was a unique intersection between their input sets, the players now have to send enough information to determine if a specific $y$ given to the referee is that unique intersection. In Section 4 we prove that solving Exam Disjointness requires at least as much information about the input as the standard Disjointness problem.
This communication model turns out to be very useful for reducing a communication problem to a one-pass streaming algorithm problem. Intuitively, only the suffix a of the stream would depend on the referee’s input $y$, which would in turn mean that the algorithm processing the rest of the stream has to learn enough information to answer any possible question of the referee that might appear in the exam. In Section 5.1, we use the Exam Disjointness problem to recover the existing $\Omega(1/\varepsilon^{2})$ lower bound for $F_{2}$-estimation, with a simple and short proof.
2.2 Direct Sum for Dependent Instances
A standard tool for proving lower bounds is proving a direct sum theorem; this means showing that solving several independent instances of a certain problem is as hard as solving every one of them separately. In our proof, we introduce an unintutive variant of a direct sum theorem in which several instances are dependent, and nevertheless solving them all is as hard as solving them separately.
$\ell=1$
$\ell=2$
$\ell=3$
Figure 2: The parts of the stream corresponding to the players of a $2^{\ell}$-party Exam Disjointness instance, for $\ell=1,2,3$.
As stated above, we first show that we can reduce the $t$-party Exam Disjointness problem to estimating the $F_{2}$ of a certain stream. We then introduce a distribution over input stream that encodes within it instances of $t$-party Exam Disjointness for multiple different values of $t$. See Figure 2 for an example of a stream in which Exam Disjointness instances with two, four, and eight players are encoded, using different parts of the stream to encode each player; while at the same level all players get a disjoint part of the stream — players at different levels occasionally intersect. We would describe a construction containing roughly $\log\left(\varepsilon^{2}n\right)$ such levels, such that solving the Exam Disjointness instance corresponding to a single level requires $\Omega(1/\varepsilon^{2})$ bits of space. If instances of different levels were disjoint, a direct sum theorem would yield our desired $\Omega\left(\log\left(\varepsilon^{2}n\right)/\varepsilon^{2}\right)$ lower bound; we obtain a similar bound despite the intersection of different levels.
The core idea of this part is the following observation: Consider an average index $j$ within the stream; at each level, look at the part of the stream that encodes the last player appearing before index $j$ in that level. Although some of these players might intersect each other, usually there is a large subset of them who are pairwise disjoint. See Figure 3 for an illustration. At index $j$ we must have information about each level, as otherwise we would not be able to finally solve the instance of that level. For the levels in which the last players are pairwise disjoint, we would be able to show that the information known about each of them by $j$ is independent. This would be enough to prove a “local” direct sum for only a subset of the levels at each index $j$, and an average over all indices would then give us a bound as good as a “global” direct sum. The exact details are given in Section 5.2.
$j$
Figure 3: An index $j$ and its preceding players from several different levels.
Another way to look at the direct sum across levels is that at each level $\ell$ we get a lower bound on the amount of information the stream typically needs to store about the $\sim 2^{\ell}$ preceding stream elements. It is then shown that these lower bounds add up, as these different-scaled “pasts” are essentially disjoint. This multi-scale phenomenon is arguably the “real reason” behind the $\log n$ multiplicative factor in the lower bound. A similar phenomenon was used to show that estimating the majority of $n$ random bits requires $\Omega(\log n)$ memory [BGW20].
2.3 Improved Algorithm for The Small Error Regime
The classic algorithm of AMS gives us a bound of $O(\log n/\varepsilon^{2})$ for $(1\pm\varepsilon)$-estimating the $F_{2}$ of a stream. Evidently, this matches our $\Omega\left(\log\left(\varepsilon^{2}n\right)/\varepsilon^{2}\right)$ lower bound only when $\varepsilon$ is polynomially larger than $n^{-1/2}$. In Section 6 we introduce a slight variant of the AMS algorithm that matches our lower bound. The modification is rather simple: we randomly partition the universe of elements into roughly $1/\varepsilon^{2}$ disjoint subsets, and then run the AMS algorithm separately (and simultaneously) on the subsets of the stream contained in each of these parts. Each subset of the stream only contains around $\varepsilon^{2}n$ elements, which would result in the improved bound. A similar modification was used before in algorithms aiming to reduce the memory-probes-per-update complexity of the AMS algorithm [TZ04].
3 Mutual Information and Streaming Algorithms
We frequently use the notions of information complexity, mutual information and entropy, see [BYJKS04] for example for more background on these notions.
A one-pass streaming algorithm receives a sequence of inputs $X=\left(X_{1},\ldots,X_{n}\right)$ one by one, and eventually outputs an answer. Denote by $M_{1},\ldots,M_{n}$ the memory transcript of the streaming algorithm. That is, $M_{i}$ is the memory state of the algorithm immediately after receiving the $i$-th input $X_{i}$. We observe that $M_{i}$ is drawn from a distribution that depends solely on $M_{i-1}$ and $X_{i}$. From now on, we think of the inputs as a distribution, thus every $X_{i}$ and every $M_{i}$ is a random variable.
The following classic observation highlights the benefit of studying the mutual information between the input distribution and the memory transcript.
Observation 3.1.
The number of memory bits used by a streaming algorithm is at least
$\max_{j}\left(I\left(M_{j}\ ;\ X\right)\right).$
Proof.
It follows immediately as $I(M_{j}\ ;\ X)\leq H(M_{j})\leq\log|\text{Support}\left(M_{j}\right)|$. ∎
We next introduce a useful lemma.
Lemma 3.2.
If the inputs $X_{i}$ are mutually independent, then
$I(M\ ;\ X)=\sum_{i=1}^{n}I\left(X_{i}\ ;\ M_{i}\ |\ M_{i-1}\right).$
Proof.
Consider the following chain of equalities,
$I(X_{1},\ldots,X_{n}\ ;\ M_{1},\ldots,M_{n})$ $=\sum_{i=1}^{n}I\left(X_{1},\ldots,X_{n}\ ;\ M_{i}\ |\ M_{<i}\right)$ (1)
$=\sum_{i=1}^{n}\left(I\left(X_{i}\ ;\ M_{i}\ |\ M_{<i}\right)+I\left(X\ ;\ M_{i}\ |\ M_{<i},X_{i}\right)\right)$ (2)
$=\sum_{i=1}^{n}\left(I\left(X_{i}\ ;\ M_{i}\ |\ M_{i-1}\right)+0\right).$ (3)
Inequalities $(1)$ and $(2)$ are applications of the chain rule of mutual information. The second part of Inequality $(3)$, that is $I\left(X\ ;\ M_{i}\ |\ M_{<i},X_{i}\right)=0$, follows as $X$ is independent of the algorithm’s random coins while $M_{i}$ is fully determined by $M_{<i},X_{i}$ and these random coins. The first part of Inequality $(3)$, that is $I\left(X_{i}\ ;\ M_{i}\ |\ M_{<i}\right)=I\left(X_{i}\ ;\ M_{i}\ |\ M_{i-1}\right)$, follows as to both $X_{i}$ and $M_{i}$ conditioning on $M_{<i}$ or on $M_{i-1}$ is equivalent, this uses the independence of $X_{i}$ from $X_{<i}$. ∎
4 Exam Variant of Multi-Party Set Disjointness
At the core of our lower bound is a variant of the Multi-Party Set Disjointness problem (denoted DISJ from now on). The DISJt communication problem is defined as follows. There are $t$ players, each player $i\in[t]$ receives a set $S_{i}$ of elements from some universe $U$; We are promised that either all sets are disjoint, or otherwise there is exactly one element that appears in all of the sets and beside it all other elements are disjoint; the players need to decide in which of the two cases their input sets are. We focus on the one-way communication model in which there is a predetermined order of the players; In its turn, a player can look at its own input set and on the communication it received from the previous player, work for an unbounded amount of time, and then send a communication to the next player; The last player needs to answer whether the sets were disjoint or contained a shared element. Chakrabarti et al. [CKS03a] proved that the total length (in bits) of the message communicated by the players throughout the protocol must be at least $\Omega\left(\frac{1}{t}\sum_{i=1}^{t}|S_{i}|\right)$ if they succeed with good probability. Intuitively, this follows as otherwise the total amount of communicated information is smaller than even a single player’s set, and thus even a single set cannot be fully described throughout the protocol’s transcript. Their lower bound is optimal and holds even in a stronger communication model called the blackboard model in which each player can see the messages communicated by all previous players, rather than just the preceding one. The same lower bound holds even in the stronger model of communication in which players can talk several times in arbitrary order and not only in a one-way fashion [Gro09, Jay09].
For our needs, we introduce a slight variant of the one-way DISJt problem which we call Exam DISJt and denote by EDISJt. In this version, we still have $t$ players with input sets $S_{i}$ that can be disjoint or have a unique common intersection, but we also have an additional special player, which we call the referee. The referee receives as input a single universe element $x\in U$. The $t$ players still communicate one by one according to order, and the last player communicates a message to the referee. Then, the referee has to decide (and succeed with probability $\geq 2/3$) whether there was an intersection between the players’ sets and that intersection is exactly the referee’s input $x$. We think of this as an “exam” for the players: If they claim there was an intersection, they also need to know enough about the intersection to tell whether or not it is $x$. In this section, we extend the lower bound for DISJt and prove that EDISJt also requires as much communication.
Let $\mathcal{P}$ be a protocol that solves EDISJt, and denote by $\Pi=\Pi(X)$ its transcript on an input $X=\left(S_{1},S_{2},\ldots,S_{t},x\right)$; the transcript consists of the message communicated by each of the $t$ players during its turn to speak. Note that $\Pi$ is a random variable that depends only on $X^{\prime}:=(S_{1},\ldots S_{t})$, as the referee does not speak and thus $x$ does not affect the transcript.
We define an input distribution for the players of EDISJt. Let $m$ be a size parameter, and let $U$ be a universe of size $|U|\geq m^{4}$. Each input set $S_{i}$ is drawn randomly and independently as a subset of $U$ of size $\lceil\frac{m}{t}\rceil$. We note that $m\leq\sum_{i=1}^{t}|S_{i}|<m+t$ and that with high probability all $t$ sets are disjoint. Denote this distribution of $X^{\prime}$ by $\mu=\mu\left(t,m,U\right)$.
In this section, we prove the following lower bound on the information complexity of EDISJt on the distribution $\mu$. We also assume that $t<\frac{m}{4\ln m}$.
Theorem 4.1.
Let $\mathcal{P}$ be a protocol that solves EDISJt with success probability $\geq 2/3$ for any input with $|S_{i}|\leq\lceil\frac{m}{t}\rceil$ for $i\in[t]$. Let $X^{\prime}=(S_{1},\ldots,S_{t})\sim\mu$ and denote by $\Pi=\Pi(X^{\prime})$ the transcript of $\mathcal{P}$ on $X^{\prime}$. Then,
$I\left(X^{\prime}\ ;\ \Pi\right)=\Omega\left(\frac{m}{t}\right).$
The proof of Theorem 4.1 is similar to standard lower bound proof for multi-party set disjointness. It proceeds in two steps: first, derive an appropriate information bound for the AND function – which is the one-bit version of disjointness; second, apply a direct sum argument to obtain the lower bound statement.
4.1 AND Lower Bounds
All proofs for the hardness of the standard Disjointess problem go through a direct sum between instances of the AND problem [BYJKS04, CKS03b, Jay09]. In the ANDt communication problem, there are $t$ players and each receives a single input bit $Y_{i}\in\mathbb{F}_{2}$; the players need to compute the “and” function $\bigwedge_{i=1}^{t}Y_{i}$. Also denote by $Y=\left(Y_{1},\ldots,Y_{t}\right)$. In the DISJt problem, we may think of each player’s input as the characteristic vector of its set $S_{i}\subset U$, and denote its coordinates by $Y_{i,u}=1$ if $u\in S_{i}$ and otherwise 00. Then, the answer to DISJt is simply $\bigvee_{u\in U}\bigwedge_{i=1}^{t}Y_{i,u}$. Due to this structure, techniques introduced by [BYJKS04] show it suffices to give a lower bound to the information complexity of ANDt to imply a lower bound for DISJt. Consider the probability distribution $\nu$ over inputs to the ANDt problem that gives probability $\frac{1}{2}$ for all input bits being 00, and probability $\frac{1}{2}$ for exactly one coordinate chosen at random being $1$ and the rest 00. Let $\Pi$ be the transcript of a communication protocol that solves ANDt with probability $>2/3$, the citations above show that $I\left(Y\ ;\ \Pi\right)=\Omega\left(\frac{1}{t}\right)$. Denote by $\Pi_{t}$ the last message communicated in the protocol $\Pi$, from which the answer has to be deduced (in a one-way communication setting this is simply the message sent by the last player). In slightly more detail, they show that as long as there is a non-negligible statistical difference between the distributions of $\Pi_{t}$ when the input is $0^{t}$ and when the input is $1^{t}$ (which there must be for solving ANDt correctly), then the above information bound holds.
Lemma 4.2 (e.g., Corollary 8 of [Jay09] [^1] In [Jay09] this statement is phrased in terms of the Hellinger distance between the two distributions rather than the total variation distance, but in the constant regime the two distances are equivalent; See for example Proposition A.2 in [BYJKS04].).
If $\|\Pi_{t}(1^{t})-\Pi_{t}(0^{t})\|_{1}>0.1$, then $I_{\nu}\left(Y\ ;\ \Pi\right)=\Omega\left(\frac{1}{t}\right)$.
For our proof, we need a slight variant of Lemma 4.2. Let $p\leq\frac{1}{t}$ be some probability. We consider a distribution $\mu=\mu_{p}$ in which every input bit $Y_{i}$ is independently chosen to be $1$ with probability $p$ and otherwise 00. We prove the following adaptation of Lemma 4.2. While it is straightforward to deduce it by repeating any of the previous proofs of Lemma 4.2, we give a proof that uses them in a black-box manner.
Lemma 4.3.
Suppose $\|\Pi_{t}(1^{t})-\Pi_{t}(0^{t})\|_{1}>0.1$, then $I_{\mu}(Y;\Pi)=\Omega(p)$.
Proof.
We observe that $\mu\left(0^{t}\right)=\left(1-p\right)^{t}\geq\left(1-\frac{1}{t}\right)^{t}\geq\frac{1}{4}$, and that for every $i\in[t]$ also $\mu\left(e_{i}\right)=p\cdot\left(1-p\right)^{t-1}\geq\frac{1}{4t}$. Therefore, we may define a Boolean random variable $D$ such that: (1) $\Pr[D=1]=\Theta(p\cdot t)$; (2) $\mu|_{D=1}=\nu$. That is, conditioned on the event $D$ happening, $\mu$ becomes $\nu$. Note that $D$ is independent of $\Pi$. Using Lemma 4.2, we get
$I_{\mu}(Y;\Pi)$ $=I_{\mu}(Y,D\ ;\ \Pi)-I_{\mu}(D;\Pi|Y)=I_{\mu}(Y,D\ ;\ \Pi)=I_{\mu}(D;\Pi)+I_{\mu}(Y;\Pi|D)$
$=I_{\mu}(Y;\Pi|D)\geq\Pr[D=1]\cdot I_{\mu}(Y;\Pi|D=1)=\Pr[D=1]\cdot I_{\nu}(Y;\Pi)$
$\geq\Theta(p\cdot t)\cdot\Omega(1/t)=\Omega(p).$
[…]
4.2 Proof of Theorem 4.1
For the remainder of the proof set
$p:=\frac{m}{2t|U|}.$
Let $Y:=\{Y^{j}\}_{j=1}^{|U|}$ be $|U|$ instances of inputs to ANDt. Let $Y_{i}:=(Y^{j}_{i})_{j=1}^{|U|}$ be the $i$-th player’s input, which consists of the $i$-th player’s inputs in all of the $|U|$ instances. Given a protocol $\Pi$ that solves EDISJt, we construct a protocol $\Pi^{\prime}(Y_{1},\ldots,Y_{t})$ that we will then use to solve all of the instances of ANDt defined above:
  1. 1.
    Each player $i\in[t]$ constructs an input set $S_{i}$ as follows:
    1. (a)
      Set $S_{i}^{\prime}:=\{j\in U:~{}Y_{i}^{j}=1\}$;
    2. (b)
      If $|S_{i}^{\prime}|>\lceil\frac{m}{t}\rceil$, send ‘Fail’, and the protocol fails;
    3. (c)
      Otherwise, set $S_{i}\subset U$ by adding $\lceil\frac{m}{t}\rceil-|S_{i}^{\prime}|$ random elements from $U\setminus S_{i}^{\prime}$ to $S_{i}^{\prime}$.
  2. 2.
    Run $\Pi(S_{1},\ldots,S_{t})$ and return the message sent to the referee as the output $\Pi^{\prime}_{t}$.
We first observe that if every instance $Y^{j}$ is drawn from $\mu_{p}$ then w.h.p no player will ‘Fail’.
Lemma 4.4.
When every $Y^{j}$ is independently distributed as $\mu_{p}$, then the probability that any $S_{i}$ is larger than $m/t$ is at most $te^{-{m}/({2t})}<\frac{1}{4m\ln m}$.
Proof.
The size of $S_{i}$ is distributed as $\text{Binomial}(|U|,\frac{m}{2t|U|})$ and therefore a Chernoff bound gives that
$\Pr\left(|S_{i}|>\frac{m}{t}\right)\leq e^{-{m}/({2t})}.$
We then take a union bound over the $t$ indices. ∎
Then, we also observe that the protocol’s output $\Pi^{\prime}_{t}$ can be used to answer each of the ANDt instances.
Lemma 4.5.
[…]
Proof.
[…]
For any $j\in U$, Lemma 4.5 implies that $\Pi^{\prime}_{t}$ conditioned on $Y^{j}=0^{t}$ must be at statistical distance $>1/3-o(1)$ from $\Pi^{\prime}_{t}$ conditioned on $Y^{j}=1^{t}$. Therefore, by Lemma  4.3 we have that when $Y^{j}\sim\mu_{p}$ (like all other instances $Y^{u}$) then
$I(Y^{j};\Pi^{\prime})=\Omega(p)=\Omega\left(\frac{m}{t|U|}\right).$
Thus if every $Y^{j}\sim\mu_{p}$ independently, we also have
$I(Y;\Pi^{\prime})=\sum_{j\in U}I(Y^{j};\Pi^{\prime}|Y^{<j})=\sum_{j\in U}I(Y^{j}\ ;\ \Pi^{\prime},Y^{<j})\geq\sum_{j\in U}I(Y^{j};\Pi^{\prime})=\Omega\left(\frac{m}{t}\right).$
Let $F$ denote the Boolean random variable such that $F=1$ if $\Pi^{\prime}$ fails (that is, any player outputs ‘Fail’). We observe that conditioned on $F=0$, the distribution of $X^{\prime}:=(S_{1},\ldots,S_{t})$, the input to the portocol $\Pi$, is exactly $\mu=\mu\left(t,m,U\right)$. Since $\Pi$ runs on the input $X^{\prime}$ when we do not fail, then $I(Y;\Pi^{\prime}\ |\ X^{\prime},F=0)=0$ as there is no additional information about $Y$ seen by the protocol besides $X^{\prime}$. We therefore have,
$I_{\mu}(\Pi;X^{\prime})$ $=I(\Pi^{\prime};X^{\prime}|F=0)=I(\Pi^{\prime};X^{\prime},Y|F=0)-I(\Pi^{\prime};Y|X^{\prime};F=0)$
$=I(\Pi^{\prime};X^{\prime},Y|F=0)\geq I(\Pi^{\prime};Y|F=0)\geq I(\Pi^{\prime};Y)-H(F)-\Pr[F=1]\cdot I(\Pi^{\prime};Y|F=1)$
$\geq\Omega\left(\frac{m}{t}\right)-1-t\exp\left(-\frac{m}{2t}\right)\cdot m\log|U|=\Omega\left(\frac{m}{t}\right).$
The last inequality holds since even in the failure event, each player who transmits anything but ‘Fail’ only holds at most $m/t$ elements, bounding the total entropy of their inputs by $m\log|U|$.
5 Lower Bound
We begin with Section 5.1, in which we show how to reduce an instance of EDISJ to the problem of $F_{2}$ estimation on a stream. This reduction already gives a simple proof of Woodruff’s $\Omega(1/\varepsilon^{2})$ lower bound [Woo04] in the entire range of $\varepsilon=\Omega(1/\sqrt{n})$. Then, in Section 5.2 we show that we can pack $\Theta\left(\log\left(\varepsilon^{2}n\right)\right)$ instances of EDISJ into a single stream of length $n$. Solving each single instance would require $\Omega\left(\frac{1}{\varepsilon^{2}}\right)$ space; We prove that solving all $\Theta\left(\log\left(\varepsilon^{2}n\right)\right)$ instances requires $\Omega\left(\log\left(\varepsilon^{2}n\right)/\varepsilon^{2}\right)$ memory bits, as much as it would take to solve each of them independently. Crucially, our instances are not going to be independent of each other, but to share certain stream elements. Finally, we prove the following, which is the main theorem of this paper.
Theorem 5.1.
Let $\mathcal{A}$ be a streaming algorithm that gives an $\left(1\pm\varepsilon\right)$ approximation to the $F_{2}$ of its input stream and succeeds with probability $\geq 2/3$, for some $\varepsilon=\Omega(1/\sqrt{n})$. Then, the space used by $\mathcal{A}$ is $\Omega\left(\log\left(\varepsilon^{2}n\right)/\varepsilon^{2}\right)$.
In Section 5.3 we extend the found to $F_{p}$ estimation for $p\in(1,2]$.
5.1 Reducing EDISJ to $F_{2}$
Fix some $n$$\varepsilon>\frac{2}{\sqrt{n}}$ and $|U|>n^{3}$. Denote by $t:=\lfloor\varepsilon\sqrt{n}\rfloor\geq 2$, for simplicity we assume that $t$ divides $n$. We reduce EDISJt to estimating $F_{2}$ on a stream of length $\left(1\pm o\left(1\right)\right)n$ up to error $\varepsilon$. Let $X=(S_{1},\ldots,S_{t},x)$ be a valid input to EDISJt, for simplicity we also assume that $|S_{i}|=\frac{n}{t}$ for every $i\in[t]$ (otherwise we may pad the sets). We define a stream $s(X)$ with elements from $U$ that encodes $X$ as follows: First, we write down the elements of $S_{1}$ in arbitrary order, afterwards we write down the elements of $S_{2}$, then of $S_{3}$ and so on until those of $S_{t}$; Finally, we write down $x$ repeatedly $k:=\lceil\frac{t}{\varepsilon}\rceil=\Theta\left(\sqrt{n}\right)$ times. We note that the length of $s(X)$ is $\left(1+O\left(\frac{1}{\sqrt{n}}\right)\right)n$, and that the stream is naturally partitioned to $t+1$ parts such that each part depends on the input of a single player (where the $(t+1)$-th ‘player’ is the referee).
Lemma 5.2.
[…]
Proof.
[…]
$2tk-2k=2k(t-1)\geq kt\geq\frac{t^{2}}{\varepsilon}\geq\varepsilon n.$
[…]
$I\left(X^{\prime}\ ;\ \Pi\right)=\Omega\left(\frac{n}{t}\right)=\Omega\left(\frac{\sqrt{n}}{\varepsilon}\right).$
Corollary 5.3.
For any $\varepsilon=\Omega(1/\sqrt{n})$, approximating $F_{2}$ up to multiplicative error $(1\pm\varepsilon)$ requires $\Omega(1/\varepsilon^{2})$ bits of space.
Proof.
The length of $\Pi=\left(M_{1},\ldots,M_{t}\right)$ is at most $t$ times the size $M$ of the space used by the streaming algorithm $\mathcal{A}$ in bits. Thus,
$\Omega\left(\frac{\sqrt{n}}{\varepsilon}\right)\leq I\left(X^{\prime}\ ;\ \Pi\right)\leq H(\Pi)\leq tM,$
and therefore
$M\geq\Omega\left(\frac{\sqrt{n}}{\varepsilon t}\right)=\Omega\left(\frac{1}{\varepsilon^{2}}\right).$
For the consecutive parts of our lower bound, we need a slightly more flexible reduction from EDISJ to $F_{2}$. In particular, we show that with a slight modification the above reduction yields the same lower bound even if we reduce from EDISJt for any $2\leq t\leq\varepsilon\sqrt{n}$ rather than exactly $t=\lfloor\varepsilon\sqrt{n}\rfloor$. Fix any such $t$. Denote by $d:=\lfloor\frac{\varepsilon^{2}n}{t^{2}}\rfloor\geq 1$. For simplicity, we assume that $d\cdot t$ divides $n$. We still use the universe $U$ for the stream elements, but our EDISJt instances are defined over the larger universe $U^{d}$ of $d$-tuples of elements from $U$. We add the promise that we only need to solve instances of EDISJt in which no two distinct elements of $U^{d}$ appearing in the input intersect each other. We consider instances of EDISJt in which the set sizes are $|S_{i}|=\frac{n}{dt}$. For an instance $X=\left(S_{1},\ldots,S_{t},x\right)$ the stream $s(X)$ is constructed almost identically to the previous construction, except that when we write an element of $U^{d}$ on the stream we write replace it with the $d$ elements of $U$ it consists of. We still repeat the referee’s element $x\in U^{d}$ at the end of the stream for $k=\lceil\frac{t}{\varepsilon}\rceil$ times. Thus, the size of the stream corresponding to $X^{\prime}$ is still $t\cdot\frac{n}{dt}\cdot d=n$ and its elements are still elements from $U$. The size of the suffix of the stream corresponding to the referee is $k\cdot d=O\left(\frac{td}{\varepsilon}\right)=O\left(\frac{t\varepsilon^{2}n}{\varepsilon t^{2}}\right)=O\left(\frac{\varepsilon n}{t}\right)$ which is at most $O(\varepsilon n)$. We next observe that Lemma 5.2 still holds: In both the ‘Yes’ and ‘No’ cases, we have $F_{2}\left(s\left(X\right)\right)\leq d\cdot\left(\frac{n}{d}+t^{2}-t+k^{2}+2tk\right)=O\left(n\right)$. Furthermore, the gap between the ‘Yes’ case and the largest possible ‘No’ case is at least $d\cdot\left(2tk-2k\right)\geq\Omega(dtk)\geq\Omega\left(\varepsilon n\right)$. We conclude as before that for $\Pi=\left(M_{1},\ldots,M_{t}\right)$ and input distribution $X^{\prime}\sim\mu\left(t,\frac{n}{d},U^{d}\right)$ we have
$I\left(X^{\prime}\ ;\ \Pi\right)=\Omega\left(\frac{n/d}{t}\right)=\Omega\left(\frac{t}{\varepsilon^{2}}\right).$
Remark 5.4.
This modification implies that the $\Omega(1/\varepsilon^{2})$ lower bound follows even from the Exam version of the $2$-party DISJ problem.
5.2 Packing many EDISJ instances into a single $F_{2}$ instance
We now construct a stream of length $n$ that encodes within it many instances of EDISJ. Fix some $n,\varepsilon>\frac{4}{\sqrt{n}}$ and $|U|>n^{3}$. For simplicity, assume that $\varepsilon^{-}1$ is a power of two and that $n$ is a power of four (otherwise we may round to the nearest ones); In particular, $\varepsilon\sqrt{n}$ is also a power of two and divides $n$.
Let $\ell\in[1,\log\left(\varepsilon\sqrt{n}\right)-2]$ be a level index. We construct a stream of length $n$ that encodes an instance of EDISJt for $t=2^{\ell}$, similar to that of Section 5.1. Denote by $d:=\frac{\varepsilon^{2}n}{4t^{2}}$. Denote by $X^{\prime}=\left(S_{1},\ldots,S_{t}\right),X=\left(X^{\prime},x\right)$ a valid instance to EDISJt with set sizes $|S_{i}|=\frac{n}{4dt}$ and universe $U^{d}$. As before, the only valid inputs are those in which all sets $S_{i}\subset U^{d}$ are either disjoint or have a unique intersection, and furthermore every distinct elements of $U^{d}$ appearing in the input do not intersect each other (i.e., each element of $U$ is used in at most one element of $U^{d}$ appearing in the input sets).
We start by partitioning the stream of length $n$ into named parts. We partition the stream into $2^{\ell+2}$ consecutive blocks of size $\frac{n}{2^{\ell+2}}$ each. We call each such block a bucket; The buckets of indices divisible by four are called active buckets and the buckets of other indices are called inactive buckets, denote by $B^{(\ell)}_{1},B^{(\ell)}_{2},\ldots,B^{(\ell)}_{2^{\ell}}$ the $t$ active buckets, each of them is a consecutive subsequence of the stream of length ${n}/{2^{\ell+2}}={n}/{4t}$. Within each active bucket, we again partition the stream into consecutive blocks of size $d$ each; We call each such sub-block a super-element. Every active bucket contains $\frac{n}{2^{\ell+2}\cdot d}=\frac{n}{4dt}$ super-elements. We refer the reader to Figure 2 for an illustration of the active buckets in several levels.
Reduction from DISJt:
We follow the reduction of Section 5.1 in our current terminology. Given a valid input $X$, we put the set $S_{i}$ for $i\in[t]$ in place of the active bucket $B^{(\ell)}_{i}$ in the stream, where each element $x\in S_{i}\subset U^{d}$ fills an entire super-element within the active bucket. Every element of the stream not within an active bucket is drawn uniformly and independently from $U$. We append to the described stream a suffix of length $d\cdot k$, for $k=\frac{t}{\varepsilon}$, which we fill with $k$ repetitions of $x$ written as a super-element each time.
[…]
Denote by $\xi$ the distribution of a stream of length $n$ where each element is drawn uniformly and independently from $U$. We observe that when $X^{\prime}\sim\mu\left(t,\frac{n}{4d},U^{d}\right)$, the resulting stream without the referee’s suffix $S(X^{\prime})$ is exactly distributed according to $\xi$. Thus, from Theorem 4.1 we conclude the following.
Corollary 5.5.
For indices $j_{1}<j_{2}<\ldots<j_{t}$ as above, we have
$I\left(\ \left(B^{(\ell)}_{1},B^{(\ell)}_{2},\ldots,B^{(\ell)}_{t}\right)\ ;\ \left(M_{j_{1}},\ldots,M_{j_{t}}\right)\ \right)=\Omega\left(\frac{t}{\varepsilon^{2}}\right),$
where the memory states of the algorithm $\mathcal{A}$ and the active buckets are with respect to a stream drawn from $\xi$.
Information measure for a single level $\ell$:
Following the above, we define an information measure that encapsulates how much information about the active buckets of level $\ell$ is given by the memory of the algorithm in an average index. From now on, we only consider the input stream distribution $\xi$ for $\mathcal{A}$, and denote the stream itself by $X=\left(X_{1},\ldots,X_{n}\right)\sim\xi$.
Definition 5.6.
For level $\ell$, denote by
${I}_{\ell}:=\sum_{j=1}^{n}I\left(X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}\ ;\ {M}_{j}\ \mid\ M_{\left(j-\frac{n}{2^{\ell}}\right)}\right).$
We next give a lower bound for $I_{\ell}$, based on Corollary 5.5.
Lemma 5.7.
$I_{\ell}>\Omega\left(\frac{n}{\varepsilon^{2}}\right)$.
Proof.
Let $j_{1}$ be any index in the range $\left(\frac{3}{4}\cdot\frac{n}{2^{\ell}},\frac{n}{2^{\ell}}\right)$. For $1<i\leq 2^{\ell}$, denote by $j_{i}:=j_{i-1}+\frac{n}{2^{\ell}}$. We observe that each $j_{i}$ is between the end of the $i$-th active bucket $B^{(\ell)}_{i}$ and the beginning of the next, so by Corollary 5.5 we have
$I\left(\ \left(B^{(\ell)}_{1},B^{(\ell)}_{2},\ldots,B^{(\ell)}_{2^{\ell}}\right)\ ;\ \left(M_{j_{1}},\ldots,M_{j_{2^{\ell}}}\right)\ \right)=\Omega\left(\frac{2^{\ell}}{\varepsilon^{2}}\right).$
By applying Lemma 3.2 this also implies
$\sum_{i=1}^{2^{\ell}}I\left(B_{i}^{(\ell)}\ ;\ M_{j_{i}}\ \mid\ M_{j_{(i-1)}}\right)>\Omega\left(2^{\ell}\cdot\frac{1}{\varepsilon^{2}}\right),$
where we abuse notation and treat $M_{j_{0}}$ as an empty conditioning. Furthermore, each active bucket $B_{i}^{(\ell)}$ is fully contained within $X_{\left(j_{i}-\frac{n}{2^{\ell}},\ j_{i}-\frac{n}{2^{\ell+1}}\right]}$; Hence,
$I\left(X_{\left(j_{i}-\frac{n}{2^{\ell}},\ j_{i}-\frac{n}{2^{\ell+1}}\right]}\ ;\ M_{j_{i}}\ \mid\ M_{j_{i-1}}\right)\geq I\left(B_{i}^{(\ell)}\ ;\ M_{j_{i}}\ \mid\ M_{j_{i-1}}\right).$
The sequences $j_{1},j_{2},\ldots,j_{2^{\ell}}$ corresponding to each different $j_{1}\in\left(\frac{3}{4}\cdot\frac{n}{2^{\ell}},\frac{n}{2^{\ell}}\right)$ are disjoint, and there are $\Omega\left(\frac{n}{2^{\ell}}\right)$ of those. We thus conclude that
$I_{\ell}\geq\Omega\left(\frac{n}{2^{\ell}}\right)\cdot\Omega\left(2^{\ell}\frac{1}{\varepsilon^{2}}\right)=\Omega\left(n\cdot\frac{1}{\varepsilon^{2}}\right).$
Combining multiple levels:
Crucially, while the analysis leading to the lower bound of $I_{\ell}$ depended on the specific level $\ell$, the input stream’s distribution is the same across all levels. Our final component is a type of direct-sum over all $\Theta\left(\log\left(\varepsilon^{2}n\right)\right)$ information measures $I_{\ell}$.
Lemma 5.8.
For any index $j\in[n]$, we have
$I\left(X_{1},X_{2},\ldots,X_{j-1}\ ;\ {M}_{j}\right)\geq\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}I\left(X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}\ ;\ {M}_{j}\ \mid\ M_{\left(j-\frac{n}{2^{\ell}}\right)}\right).$
Proof.
The subsets $X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}$ of $X_{1},\ldots,X_{j-1}$ are disjoint, and we may thus use the chain law of mutual information to get
$I\left(X_{1},X_{2},\ldots,X_{j-1}\ ;\ {M}_{j}\right)\geq\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}I\left(X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}\ ;\ {M}_{j}\ \mid\ X_{\leq\left(j-\frac{n}{2^{\ell}}\right)}\right).$
We conclude the proof by observing that
$I\left(X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}\ ;\ {M}_{j}\ \mid\ X_{\leq\left(j-\frac{n}{2^{\ell}}\right)}\right)=I\left(X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}\ ;\ {M}_{j}\ \mid\ M_{\left(j-\frac{n}{2^{\ell}}\right)}\right),$
as to both $X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}$ and ${M}_{j}$ the conditioning on either $X_{\leq\left(j-\frac{n}{2^{\ell}}\right)}$ or $M_{\left(j-\frac{n}{2^{\ell}}\right)}$ is equivalent. ∎
We next define a global measure of information between the algorithm’s memory and the inputs, then bound it using the single-level information measures.
Definition 5.9.
Denote by
$\bar{I}:=\frac{1}{n}\sum_{j=1}^{n}I\left(X_{<j}\ ;\ M_{j}\right).$
Lemma 5.10.
$\bar{I}\geq\frac{1}{n}\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}I_{\ell}$.
Proof.
Using Lemma 5.8 we have
$\sum_{j=1}^{n}I\left(X_{<j}\ ;\ {M}_{j}\right)$ $=\sum_{j=1}^{n}\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}I\left(X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}\ ;\ {M}_{j}\ \mid\ M_{\left(j-\frac{n}{2^{\ell}}\right)}\right)$
$\geq\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}\sum_{j=1}^{n}I\left(X_{\left(j-\frac{n}{2^{\ell}},\ j-\frac{n}{2^{\ell+1}}\right]}\ ;\ {M}_{j}\ \mid\ M_{\left(j-\frac{n}{2^{\ell}}\right)}\right)$
$=\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}I_{\ell}.$
Proof of Theorem 5.1.
By Lemma 5.10 and Lemma 5.7 we have
$\bar{I}\geq\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}\frac{1}{n}I_{\ell}\geq\sum_{\ell=1}^{\log\left(\varepsilon\sqrt{n}\right)-2}\Omega\left(\frac{1}{\varepsilon^{2}}\right)\geq\Omega\left(\frac{\log\left(\varepsilon^{2}n\right)}{\varepsilon^{2}}\right).$
We conclude the proof by using Observation 3.1. ∎
5.3 Generalization to $p\in(1,2]$
The lower bound presented in this section naturally translates to the problem of estimating the fractional $p$-th norm (that is, $\sum_{i}f_{i}^{p}$ for the frequency vector $\{f_{i}\}$ or equivalently its $L_{p}$ norm), for $p\in(1,2]$.
Theorem 5.11.
Fix $p\in(1,2]$. Let $\mathcal{A}$ be a streaming algorithm that gives an $\left(1\pm\varepsilon\right)$ approximation to the $F_{p}$ of its input stream, for some $\varepsilon\in\left(\Omega\left(n^{-1/p}\right),{4^{-1/(p-1)}}\right)$, and succeeds with probability $\geq 2/3$. Then, the space used by $\mathcal{A}$ is $\Omega\left(\log\left(\varepsilon^{1/p}n\right)/\varepsilon^{2}\right)$.
We only give a short sketch of this generalization, as it only requires setting the parameters of the proof for $p=2$ differently and the rest of the proof remains as-is.
We consider level indices in the range $\ell\in[1,\log\left(\varepsilon n^{1/p}\right)-2]$, all construction parameters remain as before besides $d$ which we set to $d=(\varepsilon^{p}n)/(4t^{p})$ instead. In particular, the places of the active buckets in each level remain as before. The appropriate parallel of Lemma 5.2 still holds, as the $F_{p}$ of the stream is in all cases dominated by $O\left(n+d\cdot k^{p}\right)=O(n)$ as
$d\cdot k^{p}=\left(\frac{\varepsilon^{p}n}{4t^{p}}\right)\cdot\left(\frac{t}{\varepsilon}\right)^{p}\leq n;$
Furthermore, the gap between the ‘Yes’ and ‘No’ cases is at least
$d\cdot\left(\left(t+k\right)^{p}-t^{p}-\left(k+1\right)^{p}\right)$ $=\left(\frac{\varepsilon^{p}n}{4t^{p}}\right)\cdot\left(\left(t+\frac{t}{\varepsilon}\right)^{p}-t^{p}-\left(\frac{t}{\varepsilon}+1\right)^{p}\right)$
$=\frac{1}{4}n\cdot\left(\left(1+\varepsilon\right)^{p}-\varepsilon^{p}-\left(1+\frac{1}{t}\varepsilon\right)^{p}\right)$
$\geq\frac{1}{4}n\cdot\left(\left(1+\varepsilon\right)^{p}-\left(1+\frac{1}{2}\varepsilon\right)^{p}-\varepsilon^{p}\right)$
$\geq\frac{1}{4}n\cdot\left(\frac{p}{2}\varepsilon-\varepsilon^{p}\right)=\Omega(\varepsilon n),$
where we used that $\left(1+x\right)^{p}-\left(1+\frac{x}{2}\right)^{p}\geq\frac{p}{2}x$ for every $x\geq 0,p>1$, and also that $\frac{p}{2}\varepsilon-\varepsilon^{p}\geq\frac{p}{4}\varepsilon$ when $\varepsilon\leq 4^{-1/(p-1)}$.
6 Upper Bound
In this section we give a new tight upper bound on the problem of $F_{2}$ estimation. The upper bound is similar to the original AMS construction [AMS96], with the modification that the random vectors onto which we calculate the projections have (random) disjoint support. To simplify the presentation we assume that the algorithm has access to public randomness – similarly to prior works this assumption can be relaxed by replacing true randomness with $k$-wise independence for a constant $k$. The same algorithm with a different choice of parameters appeared in [TZ04] (see also [CCFC02]), in the context of improving the update time of AMS. We include a full analysis here for completeness.
Algorithm 1 Partition-based $F_{2}$ algorithm
$P\leftarrow\lceil 4/\varepsilon^{2}\rceil+1$;
Let $H:U\rightarrow[P]$ and $\gamma:U\rightarrow\{-1,1\}$ be two random hash functions;
Initialize an array $A[1..P]$ of integer counters to 00;
for elements $x$ in the stream do
     $A[H(x)]\leftarrow A[H(x)]+\gamma(x)$
Output $A=\sum_{i=1}^{P}A[i]^{2}$;
6.1 Correctness analysis of Algorithm 1
For the analysis, let $A$ be the random variable representing the algorithm’s output, and $A_{i}$ the random variable representing the $i$-th cell of the array.
Suppose the stream of length $n$ consists of elements $\{x_{j}\in U\}_{j=1}^{n}$ with frequencies $\{f_{j}\geq 0\}_{j=1}^{n}$ [^2] We allow some frequencies to be 00 so that the counter on $j$ goes from $1$ to $n$.
Let $\gamma_{j}:=\gamma(x_{j})$, and let $B_{ji}\in\{0,1\}$ be the indicator random variable specifying whether $x_{j}$ got mapped to $A[i]$:
$B_{ji}:=\mathbf{1}_{H(x_{j})=i}.$
Then we have
$A_{i}=\sum_{j=1}^{n}f_{j}\cdot B_{ji}\cdot\gamma_{j};~{}~{}~{}A=\sum_{i=1}^{P}A_{i}^{2}$
Next, we compute $A$’s expectation and variance over the choices of the $H$ and $\gamma$ hash functions.
Observe that for each $i$,
$\mathsf{E}[A_{i}^{2}]$ $=\mathsf{E}\left[\sum_{j\in[n]}f_{j}^{2}\cdot B_{ji}^{2}\cdot\gamma_{j}^{2}+\sum_{j_{1}\neq j_{2}\in[n]}f_{j_{1}}f_{j_{2}}\cdot B_{j_{1}i}B_{j_{2}i}\cdot\gamma_{j_{1}}\gamma_{j_{2}}\right]$
$=\mathsf{E}\left[\sum_{j\in[n]}f_{j}^{2}\cdot B_{ji}\right]$
$=\frac{1}{P}\cdot\sum_{j\in[n]}f_{j}^{2}$
And for each $i_{1}\neq i_{2}$:
$\mathsf{E}[A_{i_{1}}\cdot A_{i_{2}}]$ $=\mathsf{E}\left[\sum_{j\in[n]}f_{j}^{2}\cdot B_{ji_{1}}\cdot B_{ji_{2}}\cdot\gamma_{j}^{2}+\sum_{j_{1}\neq j_{2}\in[n]}f_{j_{1}}f_{j_{2}}\cdot B_{j_{1}i_{1}}B_{j_{2}i_{2}}\cdot\gamma_{j_{1}}\gamma_{j_{2}}\right]$
$=0,$
since for all $j$, $B_{ji_{1}}\cdot B_{ji_{2}}=0$.
Therefore,
$E[A]=\sum_{i=1}^{P}E[A_{i}^{2}]=\sum_{j\in[n]}f_{j}^{2}=F_{2},$
and $A$ is an unbiased estimator of the $F_{2}$ moment.
Next, we will calculate the variance $\mathsf{Var}[A]$. Let $i_{1}\leq i_{2}\leq i_{3}\leq i_{4}\in[P]$ be any four (not necessarily distinct) elements. We have
$\mathsf{E}[A_{i_{1}}\cdot A_{i_{2}}\cdot A_{i_{3}}\cdot A_{i_{4}}]$ $=\sum_{j_{1},j_{2},j_{3},j_{4}\in[n]}f_{j_{1}}f_{j_{2}}f_{j_{3}}f_{j_{4}}\cdot\mathsf{E}[B_{j_{1}i_{1}}B_{j_{2}i_{2}}B_{j_{3}i_{3}}B_{j_{4}i_{4}}]\cdot\mathsf{E}[\gamma_{j_{1}}\gamma_{j_{2}}\gamma_{j_{3}}\gamma_{j_{4}}]$
This equals to 00 unless $i_{1}=i_{2}$ and $i_{3}=i_{4}$: for example, if $i_{1}<i_{2}\leq i_{3}\leq i_{4}$, then when $j_{1}\in\{j_{2},j_{3},j_{4}\}$ we have $B_{j_{1}i_{1}}B_{j_{2}i_{2}}B_{j_{3}i_{3}}B_{j_{4}i_{4}}=0$, and when $j_{1}\notin\{j_{2},j_{3},j_{4}\}$, we have $\mathsf{E}[\gamma_{j_{1}}\gamma_{j_{2}}\gamma_{j_{3}}\gamma_{j_{4}}]=0$.
Therefore, we only need to consider terms of the form $\mathsf{E}[A_{i_{1}}^{2}\cdot A_{i_{2}}^{2}]$.
If $i_{1}<i_{2}$, we have
$\mathsf{E}[A_{i_{1}}^{2}\cdot A_{i_{2}}^{2}]$ $=\sum_{j_{1},j_{2},j_{3},j_{4}\in[n]}f_{j_{1}}f_{j_{2}}f_{j_{3}}f_{j_{4}}\cdot\mathsf{E}[B_{j_{1}i_{1}}B_{j_{2}i_{1}}B_{j_{3}i_{2}}B_{j_{4}i_{2}}]\cdot\mathsf{E}[\gamma_{j_{1}}\gamma_{j_{2}}\gamma_{j_{3}}\gamma_{j_{4}}]$
$=\sum_{j_{1}\neq j_{3}\in[n]}f_{j_{1}}^{2}f_{j_{3}}^{2}\cdot\mathsf{E}[B_{j_{1}i_{1}}B_{j_{3}i_{2}}]\cdot\mathsf{E}[\gamma_{j_{1}}^{2}\gamma_{j_{3}}^{2}]$
$=\sum_{j_{1}\neq j_{3}\in[n]}f_{j_{1}}^{2}f_{j_{3}}^{2}/P^{2}$
$=\frac{1}{P^{2}}\cdot\left(F_{2}^{2}-\sum_{j}f_{j}^{4}\right)$
$=\frac{1}{P^{2}}\cdot\left(F_{2}^{2}-F_{4}\right)$
If $i_{1}=i_{2}$, we have
$\mathsf{E}[A_{i}^{4}]$ $=\sum_{j_{1},j_{2},j_{3},j_{4}\in[n]}f_{j_{1}}f_{j_{2}}f_{j_{3}}f_{j_{4}}\cdot\mathsf{E}[B_{j_{1}i}B_{j_{2}i}B_{j_{3}i}B_{j_{4}i}]\cdot\mathsf{E}[\gamma_{j_{1}}\gamma_{j_{2}}\gamma_{j_{3}}\gamma_{j_{4}}]$
$=6\cdot\sum_{j_{5}<j_{6}\in[n]}f_{j_{5}}^{2}f_{j_{6}}^{2}\cdot\mathsf{E}[B_{j_{5}i}B_{j_{6}i}]+\sum_{j\in[n]}f_{j}^{4}\cdot\mathsf{E}[B_{ji}]$
$=\frac{3}{P^{2}}\cdot\left(F_{2}^{2}-\sum_{j}f_{j}^{4}\right)+\frac{1}{P}\cdot\sum_{j}f_{j}^{4}$
$=\frac{3}{P^{2}}\cdot\left(F_{2}^{2}-F_{4}\right)+\frac{1}{P}\cdot F_{4}$
We are now ready to compute $\mathsf{Var}[A]$:
$\mathsf{Var}[A]$ $=\mathsf{E}[A^{2}]-F_{2}^{2}$
$=\sum_{i_{1}\neq i_{2}\in[P]}\mathsf{E}[A_{i_{1}}^{2}\cdot A_{i_{2}}^{2}]+\sum_{i\in[P]}\mathsf{E}[A_{i}^{4}]-F_{2}^{2}$
$=P\cdot(P-1)\cdot\frac{1}{P^{2}}\cdot(F_{2}^{2}-F_{4})+P\cdot\left(\frac{3}{P^{2}}\cdot\left(F_{2}^{2}-F_{4}\right)+\frac{1}{P}\cdot F_{4}\right)-F_{2}^{2}$
$=\frac{2}{P}\cdot F_{2}^{2}-\frac{2}{P}\cdot F_{4}$
$<\frac{2}{P}\cdot F_{2}^{2}$
Theorem 6.1.
The expected relative $\ell_{2}$-error of Algorithm 1 is bounded by $\varepsilon$:
$\mathsf{E}\left[\left(\frac{A-F_{2}}{F_{2}}\right)^{2}\right]<\varepsilon^{2}$ (1)
Proof.
Since $\mathsf{E}[A]=F_{2}$, we have
$\mathsf{E}[(A-F_{2})^{2}]=\mathsf{Var}[A]<\frac{2}{P}\cdot F_{2}^{2}<\varepsilon^{2}\cdot F_{2}^{2},$
implying (1). ∎
Memory analysis of Algorithm 1
In the word model, or if each element $A[i]$ of the array is stored using $O(\log n)$ bits, the memory cost of the algorithm is $O(P\log n)=O(\varepsilon^{-2}\cdot\log n)$ – matching the memory cost of [AMS96]. We observe that in some regimes – when $\log(\varepsilon^{2}n)\ll\log n$ – Algorithm 1 requires asymptotically less memory, matching our lower bounds above.
Claim 6.2.
Suppose the stream length is $n\geq\Omega(\varepsilon^{-2})$. Then Algorithm 1 can be implemented using $O(P\log(n/P))=O(\varepsilon^{-2}\cdot\log(\varepsilon^{2}n))$ memory.
Proof.
Note that the state of the memory consists of $P$ integers satisfying $\sum_{i}|A[i]|\leq n$. Each integer $A[i]$ can be stored using a prefix-free code that only requires $2\log(|A[i]|+1)+O(1)$ bits [^3] In fact, $\log|A[i]|+O(\log\log|A[i]|)$ suffices.
The total memory cost of concatenating $P$ such prefix-free encodings is therefore at bounded by
$2\sum_{i\in[P]}\log(|A[i]|+1)+O(P)\leq 2P\cdot\log\left(\frac{\sum_{i\in[P]}|A[i]|}{P}+1\right)+O(P)=O(P\log(n/P)).$
Here the inequality follows from the concavity of the $\log$ function. ∎
7 Summary and Open Problems
After this work, the space complexity of $F_{p}$ estimation is understood with nearly-tight bounds for the entire range of $p\geq 0$, as illustrated in Figure 1. Nonetheless, several natural questions remain open.
For second moment estimation, we prove a tight space bound for all $\varepsilon=\Omega(1/\sqrt{n})$, which is $\Theta(n)$ when $\varepsilon=\Theta(1/\sqrt{n})$. By simply maintaining the frequency vector, an exact computation of the second moment is possible with $O\left(\min\{n\log n,|U|\}\right)$ space. A gap between $\Theta(n\log n)$ and $\Theta(n)$ space remains for the range of $0<\varepsilon<1/\sqrt{\varepsilon}$. What is the exact space complexity of $F_{2}$ estimation in the regime of very low error?
Contrasting our result with the $\tilde{O}\left(\log n+1/\varepsilon^{2}\right)$ upper bound in random-ordered streams for $p\in(1,2)$ of [BVWY18], we end up with the first proven gap between the space complexity of $F_{p}$ estimation in adversarial and random-ordered streams for any range of $p$. For $p>2$, it is known that no such gap exists [GH09]. Therefore, only the case of $p=2$ remains open: Is $F_{2}$-estimation “cheaper” in random-ordered streams?
Acknowledgments
The authors would like to deeply thank David Woodruff for several helpful conversations.
MB’s research is supported in part by the NSF Alan T. Waterman Award, Grant No. 1933331.
OZ’s research is supported in part by the Israel Science Foundation, Grant No. 1593/24.